OpenStudy (anonymous):

Find the vertex of the quadratic equation: y = 2x2 + 12x − 4. (−3, 50) (3, 50) (−3, −22) (−3, −58)

OpenStudy (anonymous):

Use the formula x = -b/2a = -12/ 4 = -3 Now we can put this values of x to find y

OpenStudy (anonymous):

-12 i knew it was -3, somethin

OpenStudy (anonymous):

12 isnt in any of the answers. sorry OS cut me off :/

OpenStudy (anonymous):

So you have x = -3 and in the equation y = 2x2 + 12x − 4. Put x = 3 to get y CAn you do that?

OpenStudy (anonymous):

x = 3 is still 3

OpenStudy (anonymous):

(−3, −22)

OpenStudy (anonymous):

yes its correct, you did that on you own..hahaha